OpenStudy (anonymous):

olve for x: −5|x + 1| = 10 x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution

OpenStudy (paki):

simple..... just do this.... −5(x + 1) = 10 and −5(x + 1) = -10... now solve....

hartnn (hartnn):

when you divide by -5 on both sides ,what do u get ?

hartnn (hartnn):

remember : |....| can never be negative!

OpenStudy (anonymous):

-5x-5=10 =-3

OpenStudy (anonymous):

so 3

hartnn (hartnn):

no... first step is to divide both sides by -5 in −5|x + 1| = 10 what do u get ?

OpenStudy (acxbox22):

no just divide by -5 on both sides those arent parentheses it is the absolute value symbol paki got u confused...

OpenStudy (anonymous):

i have no clue

OpenStudy (acxbox22):

what is 10/-5???

OpenStudy (anonymous):

-2

hartnn (hartnn):

yes, so u get |x+1| = -2 right ?

OpenStudy (anonymous):

ya

hartnn (hartnn):

is that even possible ?

OpenStudy (anonymous):

no

hartnn (hartnn):

remember : answer to absolute value sign can never be negative

hartnn (hartnn):

so whats your conclusion ?

OpenStudy (anonymous):

so its 2

hartnn (hartnn):

no, its not 2 since |x+1| can never = -2 there is no value of 'x' !

OpenStudy (anonymous):

so its no solution

hartnn (hartnn):

yessss :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can you help me with one more please

hartnn (hartnn):

welcome ^_^ and since you're new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) sure!

OpenStudy (anonymous):

thank you :)))

OpenStudy (anonymous):

-3|x + 7| = -12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

hartnn (hartnn):

divide both sides by -3

hartnn (hartnn):

always first try to isolate the absolute value sign

OpenStudy (anonymous):

4

OpenStudy (anonymous):

im really connfused with absolute values

OpenStudy (acxbox22):

now 4 is positive so go on to the next step, x+7=4 and x+4=-7 now solve for x

OpenStudy (anonymous):

-3 and -3

hartnn (hartnn):

general rule! when |a| = b we have a= b or a = -b

OpenStudy (acxbox22):

no.. x+7=4 subtract 7 on both sides x+7=-4 subtract 7 on both sides

OpenStudy (anonymous):

nvm its -3 and -11

OpenStudy (acxbox22):

good job

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

and i have one more

OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution

hartnn (hartnn):

add 4 on both sides...

OpenStudy (anonymous):

24

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

9

hartnn (hartnn):

|2x+6| = 24 now use the general rule which i gave you!

hartnn (hartnn):

when |a| = b we have a= b or a = -b when |2x+6|= 24 ....?....

OpenStudy (anonymous):

9

OpenStudy (anonymous):

is it no solution

OpenStudy (anonymous):

it has to be one of these x = 9 and x = −15 x = 9 and x = 11 No solution

hartnn (hartnn):

9 is one of the solution, whats another ?

hartnn (hartnn):

2x+6 = 24 gave you x = 9 2x+6 = -24 gives you what ?

OpenStudy (anonymous):

-15

OpenStudy (anonymous):

so its 9 and -15

hartnn (hartnn):

yes! correct :)

OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14 x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution

OpenStudy (anonymous):

is the answer to this one -32 and -4?

hartnn (hartnn):

this is easy subtract 16 from both sides what do u get ?

hartnn (hartnn):

and no, thats not the answer

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

i think its no solution

hartnn (hartnn):

can |x+2| ever be -2 ??

OpenStudy (anonymous):

no

hartnn (hartnn):

correct, so its no solution :)

OpenStudy (anonymous):

The minimum and maximum temperature for a day in Cupcake Town can be modeled by the equation below: 3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day? x = 1, x = 17 x = −7, x = 9 x = 7 x = 9 No solution

OpenStudy (anonymous):

is this one x=7 x=9

hartnn (hartnn):

ask new question in a new post :) so that others can help you too! bdw, you are correct for this :)

OpenStudy (anonymous):

ok thanks

hartnn (hartnn):

welcome ^_^